2. Use the Laplace transform to solve Y – 2y = 2 y(0) = 0, y(0) = 0
Answer #1

2 1 R = S-a - S Rule 1:- 1 {9f; (4) +2222641} = 0,2.54838 + Ag L}{U} Rule 2: L{ext] Rule3:- {{f(t)} = SZL {f4} -S610)-f(o)[ Rule jaett 5:- £?$} Ruce bir la L}} IY boer Rules:- L} F(s)} = f(t) , then L} FCS-aj} - eat fit) = eat { } F(s) Ruler: Il